# Structure of the free interfaces near triple junction singularities in harmonic maps and optimal partition problems

### v2

@@ 1. Introduction
By the works of Gromov-Schoen [GS92], Caffarelli-Lin [CL07, CL08], Tavares-Terracini [TT12] and Soave-Terracini [ST15] it is known that if
@@
By [GS92, CL08], in every dimension $d\geq 2$, the regular part $\mathrm{Reg}(u)$ is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92]) and is $(d-2)$-rectifiable ([Alp20, Dee22]). In dimension $d=2$, Conti, Terracini and Verzini showed in [CTV03] that the singular set is discrete and gave a complete description of the nodal set $\mathcal{F}(u)$ around any singular point (see also the recent paper [LMNS24] for the case of generic solutions).
@@
Results in the spirit of Theorem 1.1 exist in the context of area-minimizing currents, where the nodal set $\mathcal{F}(u)$ is replaced by the support of an area-minimizing current (mod $p=3$); in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93], in any dimension $d\geq 3$ (see also the lecture notes [Min24]); we also remark that the analysis for general co-dimension 1 area-minimizing currents mod p was completed only recently in [DHMS20], [DHM+21] and [DHM+22]. To our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$.
@@
There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on the dimension $d$ such that the following holds. For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that
@@
there exists $u\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $u=c$ on $\partial B_{1}$ and
@@
The proof of Theorem 1.2 is obtained via a contradiction argument in the spirit of Weiss [Wei99] (see also [FS16, GPS16]). The epiperimetric inequality approach to the analysis of the singular points in the context of harmonic maps is, to our knowledge, new in this field.
@@ 5. Epiperimetric inequality at points of frequency $\nicefrac{{3}}{{2}}$
Lemma 5.3.
For any $u\in H^{1}(B_{1};\Sigma_{N})$ and any $\gamma\geq 0$, we let
@@
Theorem 5.4.
There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on $d$ such that the following holds. For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $c_{n,i}\geq 0$ for all $i=1,\dots,N$,
and
$W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
Proof.

### v3

@@ 1. Introduction
By the works of Gromov-Schoen [GS92], Sun [Sun03], Caffarelli-Lin [CL07, CL08], Tavares-Terracini [TT12] and Soave-Terracini [ST15] it is known that if
@@
By [GS92, Sun03, CL08], in every dimension $d\geq 2$, the regular part $\mathrm{Reg}(u)$ is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92, Sun03] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92, Sun03]) and is $(d-2)$-rectifiable ([Alp20, Dee22]). In dimension $d=2$, Conti, Terracini and Verzini showed in [CTV03] that the singular set is discrete and gave a complete description of the nodal set $\mathcal{F}(u)$ around any singular point (see also the recent paper [LMNS24] for the case of generic solutions).
@@
Results in the spirit of Theorem 1.1 exist in the context of area-minimizing surfaces; in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93] in any dimension $d\geq 3$. The techniques from [Tay73, Tay76, Sim93] have been abundantly explored in the last decade (see for instance [DHMS20, DHM+21, DHM+22] and the references therein), but fine results as Theorem 1.1 have been obtained only for the area functional and its generalizations so far. In fact, to our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$.
@@
There exists $\delta,\varepsilon,\tau\in(0,1)$ depending only on the dimension $d$ such that the following holds. For any $N\in\mathbb{N}$ and any $\nicefrac{{3}}{{2}}$-homogeneous $c\in H^{1}(B_{1};\Sigma_{N})\cap C^{0,1}(B_{1};\mathbb{R}^{N})$ such that $c_{i}\geq 0$ for all $i=1,\dots,N$,
@@
The proof of Theorem 1.2 is obtained via a contradiction argument in the spirit of Weiss [Wei99] (see also [FS16, GPS16, Tay76, Tay73]). The epiperimetric inequality approach to the analysis of the singular points in the context of harmonic maps is, to our knowledge, new in this field.
One of the main consequences of Theorem 1.2 is the uniqueness of the ($\nicefrac{{3}}{{2}}$-homogeneous) blow-up at every point of frequency $\nicefrac{{3}}{{2}}$. We notice that the uniqueness of the tangent maps was claimed in [CL10, Section 3]. The argument in [CL10] is based on a Monneau-type formula obtained via the Weiss-energy linearization identity from [CL10, Lemma 3.2]. Unfortunately, there is a missing term in [CL10, Lemma 3.2], which compromises the conclusion in [CL10]. The correct formula is given in Lemma 5.3 and contains an extra term $\beta$. This term is linear and positive, and requires particular attention in the proof of the epiperimetric inequality. Finally, we also point out that the term $\beta$ from Lemma 5.3 is specific for our functional and contains information on the interaction between the different phases.
Our epiperimetric inequality shares some common features with the epiperimetric inequalities from [Tay76, Tay73], for instance, the concept of “linearization” and the difficulties related to the “parametrization” over the limit cone. On the other hand, our functional is of different nature, which is reflected, in particular, in the classifications of the solutions of the linearized problem in Section 4, in the Weiss-energy linearization identity from Lemma 5.3, and in the constructions of the competitors in Step 1 and Step 4 of the proof of the epiperimetric inequality in Section 5, where we face structural problems caused by the segregation constraint which is not present in [Wei99, Tay76, Tay73].
Once we have the epiperimetric inequality Theorem 1.2, the conclusion in Theorem 1.1 follows mainly from the classification of the $\nicefrac{{3}}{{2}}$-homogeneous tangent maps (Proposition 3.5), from the frequency gap from above (Lemma 6.5), and from a topological argument (Lemma 6.6) similar to [Sim93, Tay76, Tay73], which guarantees the persistence of singularities over the triple junction model.
@@
1.2. The Bishop-Friedland-Hayman’s conjecture for the min-max
Consider, for $p\in(0,+\infty]$, the family of optimal partition problems
($\text{OP}_{p}$) $\min\left\{\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\},$
where $\lambda_{1}(\omega)$ denotes the first Dirichlet eigenvalue of the spherical Laplacian in $\omega\subseteq\partial B_{1}$, and
$\left(\sum_{i=1}^{N}\lambda_{1}(\omega_{i})^{p}\right)^{\frac{1}{p}}:=\max_{i\in\{1,\dots,N\}}\lambda_{1}(\omega_{i})\quad\text{when}\quad p=+\infty.$
In the particular case of the $3$-partition, i.e. $N=3$, we have the following conjecture.
Problem 1.5 (Bishop-Friedland-Hayman-Helffer-Hoffmann-Ostenhof-Terracini).
For $N=3$, prove that the $Y$-configuration is the unique minimizer of ($\text{OP}_{p}$), for any $p\in(0,+\infty]$.
This conjecture was posed in [Bis92] in the case $p=1$, but can be generalized for any $p\in(0,+\infty]$ (see [FH76] and [HHT10]) and is related to shape optimization and other optimal partition problems [CTV03, CTV05a, BBH98, Hel10, HH10, HHT10]. 1.5 was solved for $p=+\infty$ in dimension $d=3$ by Helffer, Hoffmann-Ostenhof and Terracini in [HHT10]. In Section 3.2 we solve 1.5, for $p=+\infty$ and in any dimension $d\geq 3$, as a consequence of the classification of $\nicefrac{{3}}{{2}}$-homogeneous tangent maps (first obtained in [ST15]), see Corollary 3.6.
Organization of the paper
@@ 5. Epiperimetric inequality at points of frequency $\nicefrac{{3}}{{2}}$
In this section we prove Theorem 1.2; we start with the following general lemma, which will be useful in the construction of the competitors in Step 1 and Step 4 of the proof.
@@
The following lemma contains a general integration by parts identity for tangent maps.
@@
The next lemma is the Weiss-energy linearization identity which is a key element in the argument in the proof of Theorem 1.2.
Lemma 5.3 (Weiss-energy linearization identity).
For any $u\in H^{1}(B_{1};\Sigma_{N})$ and $\gamma\geq 0$, let
@@
We conclude this section with the following.
Proof of Theorem 1.2.
@@ Acknowledgements
Corresponding author
Bozhidar Velichkov (e-mail: bozhidar.velichkov@unipi.it)
Conflict of interests and data availability statements
The authors have no competing interests to declare. Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.